OpenStudy (loser66):

Let \(y_1:=\sqrt p\), p > 0, \(y_{n+1}=\sqrt {p +y_n}\), \(\forall n \in N\) Show \(\{y_n\}\) converges and find limit. Please, help

OpenStudy (loser66):

To show \(\{y_n\}\) converges, we need show it monotone and bounded.

OpenStudy (loser66):

\(y_2 = \sqrt{p+y_1}\\y_2-y_1=\sqrt{p+y_1}-y_1\)

OpenStudy (loser66):

@phi @ganeshie8

OpenStudy (loser66):

@hartnn @myininaya

OpenStudy (loser66):

\(p>p+\sqrt p\\\sqrt p>\sqrt{p+\sqrt p}\) \(\sqrt{p+\sqrt p}- \sqrt p<0\) replace \(\sqrt p =y_1\) we have \(\sqrt {p+ y_1}-y_1<0\) and \(y_2 =\sqrt{p+y_1}\) so that \(y_2 -y_1 <0\\y_2<y_1\) \(\{y_n\}\) decreases for n=1,2

OpenStudy (loser66):

oh, I messed up, < ,not > so that , it is increasing for n = 1,2 . I am sorry.

OpenStudy (xapproachesinfinity):

yeah i was gonna say that lol

OpenStudy (loser66):

But whatever, What I need is induction step for n = k+1

ganeshie8 (ganeshie8):

whats your upperbound ?

ganeshie8 (ganeshie8):

you need to show that the sequence is bounded first, right ?

OpenStudy (loser66):

I don't know. Now I just show it converges, after then find bound by other process. That process is .... a mess.

OpenStudy (loser66):

No, @ganeshie8 I have to prove it converges first, from then, I can assume that it is bounded and find out the bound later.

OpenStudy (loser66):

And that is lim,

ganeshie8 (ganeshie8):

{y_n} converges if it is : 1) bounded 2) monotone

OpenStudy (loser66):

Ok, if so, what should I have to do? find bounds, right? btw, I got induction step. :) hihihi Now I find bound

ganeshie8 (ganeshie8):

for bounded, show that \(\large y_n \lt L\) for all \(\large n\) for monotone, show that \(\large y_n \lt y_{n+1}\) for all \(\large n\)

ganeshie8 (ganeshie8):

let me go through your induction step again, you're trying to prove the sequence is increasing right ?

OpenStudy (loser66):

yes,

OpenStudy (loser66):

just repeat the step again with \(y_{k+2} \) and \(y_{k+1}\). hehehe.

OpenStudy (loser66):

Now, find bounded: Let x = lim \(y_n\) then x = lim \(y_{n=1}\) also ( while x --> infty\) so that \(x = \sqrt{p +x}\) square both sides \(x^2 = p+x\) \(x^2-x-p=0\) \(x= \dfrac{1\pm \sqrt{1+4p}}{2}\) since 1+4p> 1 for all p>0, we reject the minus value which leads x to negative number so that the bounds of \(\{y_n\}\) is \(\dfrac{1+\sqrt{1+4p}}{2}\)

OpenStudy (perl):

for example you want to show that sqrt( 4 + sqrt( 4 + sqrt( 4 + ... )))) converges

OpenStudy (loser66):

I don't get what you mean

OpenStudy (perl):

your question is asking the limit of square roots, infinite square rooting

OpenStudy (perl):

if we let p = 4 , for example

OpenStudy (loser66):

Nope, my question is a proving one.

OpenStudy (loser66):

After proving, find the limit.

ganeshie8 (ganeshie8):

both parts look good to me !

OpenStudy (loser66):

yyyyyyyyes! Thanks for comfirming.

OpenStudy (perl):

i was catching up *

ganeshie8 (ganeshie8):

\[\lim\limits_{n\to \infty} y_{n+1} = \lim\limits_{n\to \infty} y_{n} = L \implies L = \sqrt{p+L} \implies L = \dfrac{1+\sqrt{1+4p}}{2}\]

OpenStudy (perl):

also you can mention that it is bounded below by zero, since all the terms are positive and adding /square rooting is positive

ganeshie8 (ganeshie8):

you have proved that the sequence is increasing in your first step itself which is bit unusual as you could have used the upper bound in proving the sequence is monotone also. but your logic sounds great !

OpenStudy (loser66):

Show me your stuff, please @ganeshie8 . I need know as many method as possible.

ganeshie8 (ganeshie8):

Notice that the work for upper bound is independent of monotone part

OpenStudy (perl):

im going over your proof, {yn} is actually an increasing sequence

ganeshie8 (ganeshie8):

Having known that the sequence is bounded above by \(\large L\), we go ahead and prove that the sequence is increasing using induction

OpenStudy (perl):

where did you show the sequence is bounded above by L ?

ganeshie8 (ganeshie8):

in the earlier limit

OpenStudy (perl):

I must have missed it, i dont see an upper bound

OpenStudy (perl):